# CLIMB: CLustering-based Iterative Data Mixture Bootstrapping for Language Model Pre-training

### v1

@@ Abstract
Abstract: Pre-training datasets are typically collected from web content and lack inherent domain divisions. For instance, widely used datasets like Common Crawl do not include explicit domain labels, while manually curating labeled datasets such as The Pile is labor-intensive. Consequently, identifying an optimal pre-training data mixture remains a challenging problem, despite its significant benefits for pre-training performance. To address these challenges, we propose CLustering-based Iterative Data Mixture Bootstrapping (CLIMB), an automated framework that discovers, evaluates, and refines data mixtures in a pre-training setting. Specifically, CLIMB embeds and clusters large-scale datasets in a semantic space and then iteratively searches for optimal mixtures using a smaller proxy model and a predictor. When continuously trained on 400B tokens with this mixture, our 1B model exceeds the state-of-the-art Llama-3.2-1B by 2.0%. Moreover, we observe that optimizing for a specific domain (e.g., Social Sciences) yields a 5% improvement over random sampling. Finally, we introduce ClimbLab, a filtered 1.2-trillion-token corpus with 20 clusters as a research playground, and ClimbMix, a compact yet powerful 400-billion-token dataset designed for efficient pre-training that delivers superior performance under an equal token budget. We analyze the final data mixture, elucidating the characteristics of an optimal data mixture.
Datasets on Hugging Face: ClimbMix | ClimbLab | Homepage: CLIMB
@@ 1 Introduction
Pre-training datasets for large language models (LLMs) have scaled to trillions of tokens, typically combining large-scale web crawls with smaller, high-quality domain-specific datasets. These corpora enable the development of generalist models capable of addressing diverse tasks. However, their vast scale and heterogeneity pose challenges in balancing general knowledge with domain expertise, often leading to inefficient utilization of high-value data for specialized capabilities. Recent studies emphasize the importance of the final stage of pertaining, commonly referred to as mid-training, where models are refined on targeted, high-quality data to enhance specific capabilities. For example, [1] demonstrated that emphasizing domain-specific datasets during the final pre-training phase significantly improves performance on benchmarks such as GSM8K [2] (math), MMLU [3] (reasoning), and HumanEval [4] (coding). Similarly, OLMo 2 [5] mixes high-quality web data with curated STEM references, synthetic math datasets, and encyclopedic content for mid-training, achieving notable gains in math reasoning tasks. These findings highlight the potential of carefully curated data mixtures in mid-training for improving domain performance.
Figure 1: Pre-training a 1B model on ClimbMix shows better scaling effects than training on other datasets. We measure the average performance on 12 downstream benchmarks.
Despite the success of pre-training, optimizing data mixtures for both general and domain-specific tasks remains a challenge: (1) Large-scale datasets such as Common Crawl  [^1] https://commoncrawl.org/ offer unmatched diversity and scale but lack explicit domain labels, making it difficult to extract domain-relevant content. Filtering data often relies on general-purpose heuristics like perplexity or educational value [6], which do not necessarily capture the most informative or high-quality content for specific domains. (2) Even with curated datasets like The Pile [7] with domain annotations, selecting an optimal data mixture is non-trivial due to the complex, nonlinear relationship between dataset composition and model performance. For instance, optimizing a model for coding tasks requires not just programming-related content but also complementary knowledge from mathematics, reasoning, and security.
@@
To address these challenges, we propose CLustering-based Iterative Data Mixture Bootstrapping (CLIMB)—a novel framework for automating the search for optimal data mixtures during pre-training. CLIMB consists of three key steps: (1) embedding and clustering large-scale datasets, (2) constructing mixture-performance pairs by sampling and pruning data mixtures and training proxy models, and (3) fitting a predictor. By treating the data mixture as input features and performance metrics as target labels, we train a regression model as a predictor. This approach enables efficient, iterative refinement of data mixtures without relying solely on predefined domain labels.
@@
We demonstrate the effectiveness of CLIMB by searching the optimal data mixture in general reasoning tasks first and then extending it to specific domains (e.g., STEM, social sciences, and humanities). Using the optimal data mixture discovered by CLIMB, we train 350M and 1B models on 40B tokens, both of which surpass the previously best data mixing (Doremi and RegMix) methods by a large margin. Furthermore, when trained on a larger number of tokens (400B) with this mixture, our 1B model exceeds the state-of-the-art Llama-3.2-1B by 2.0%. We observe that optimizing for a specific domain (e.g., Social Sciences) yields a 5% improvement over random sampling. Finally, based on the insights obtained from our explorations, we further apply CLIMB to two existing datasets, Nemotron-CC [8] and smollm-corpus [9], and produce a new dataset with superior performance.
@@
    Automated Data Mixture Optimization. We propose an embedding-driven data mixing approach to automatically identify, group, mix high-quality clusters, enabling efficient domain-specific training while removing the reliance on manually predefined domain-specific data.
@@
2 Related Work
Data Mixture for LLM Pre-training. The composition of pre-training datasets are critical in determining the generalization abilities of language models  [10, 11, 12]. Typically, data mixtures like those in the Pile [7], GLaM [13], and ROOTS [14] are crafted using manually defined rules, yet these heuristics lack standardization and transferability across different settings. SlimPajama-DC [15] systematically evaluated the influence of various predefined data configurations, yielding valuable insights. More recently, learning-based approaches such as DoReMi [16] and DoGE [17] have introduced optimization techniques for domain proportions by iteratively refining training with reference and proxy models. While [18] investigated data sequencing strategies through the lens of curriculum learning, our work focuses on the simultaneous integration of diverse data domains, emphasizing a distinct aspect of pre-training. The aforementioned methods show promise, but they require the dataset to already possess clear and natural domain distinctions. By contrast, we propose a novel approach that can automatically identify approximate domains from large amounts of web data and then find the optimal data mixture automatically. In parallel work, WebOrganizer [19] proposes using classifiers to annotate web-scale data with topic and format labels. In contrast, our clustering-based approach is more straightforward, and readily scalable, and we introduce an iterative optimization method to refine the data mixture.
Data Selection for Specific Domains. Beyond optimizing the overall pre-training data mixture [20, 21, 22, 23], selecting high-quality domain-specific data [24, 25] is essential for improving model specialization during pre-training. Existing methods approach this challenge differently. DSIR [26] estimates relevance using hashed n-grams and resamples data to better match target domain distributions. CRISP [27] clusters the generalist dataset and samples these clusters according to their frequencies in the smaller specialist dataset. [28] propose to select data that nudges the pre-training distribution closer to the target distribution. Training dynamics-based selection leverages model learning behavior to guide data filtering, including S2L [29], which clusters data based on loss trajectories to prioritize domain-relevant examples, and LESS [30], which selects instruction tuning data with the highest gradient similarity to a target task. Embedding-based filtering removes redundant [31] or low-quality data, with SCIP [32] applying synthetic corruptions for filtering and heuristic pruning [33] reducing noise from overrepresented long-text clusters. [34] proposes to select the data on which model losses are predictive of downstream abilities. While these approaches improve data quality for specialized domains, they often rely on predefined domain labels or heuristics, limiting their flexibility for large-scale pre-training. In contrast, our proposed framework, CLIMB, iteratively refines domain-relevant data mixtures without requiring explicit domain labels, making it more applicable to real-world pre-training data and easier to scale.
3 CLIMB: CLustering-based Iterative Data Mixture Bootstrapping
Our work focuses on curating training data from a massive data source in an automated fashion, specifically tailored to improve the desired tasks or domains. To ensure that the filtered dataset remains relevant to the target domain while maintaining general language modeling and reasoning capabilities, our framework simplifies the data curation process through a fully autonomous iterative bootstrapping approach, eliminating the need for manual curation and reducing labor costs. As illustrated in Figure 4, we first cluster documents from the data source in an embedding space to differentiate data across domains. Next, we iteratively optimize the mixture weights using a bootstrapping process to progressively enhance the dataset’s domain relevance. Further details of the two phases are provided in Section 3.1 and 3.2, respectively.
@@ 3.1 Data Preprocessing
Embedding clustering. We then cluster the embeddings using a suitable clustering algorithm. For instance, k-means [35] can be used to group them into $K_{\text{init}}$ clusters. To ensure the clusters are as fine-grained as possible for subsequent processing, we prefer to set $K_{\text{init}}$ to a relatively large value at this stage, such as $1000$. The specific settings are detailed in Section 4.1.
@@ 3.2 Iterative Bootstrapping: Mixture Weight Search
$\min_{\alpha\in A}\,\,\ell_{val}(\alpha,\omega^{*}(\alpha))$ (1)
[…]
@@
(Sampling) (3)
$\tilde{P}^{k}=\{f_{k}(s)|s\in A\setminus S^{k}\},$
$S_{M}\subset\text{Top}_{N}(\tilde{P}^{k}),S^{k+1}=S_{M}\cup S^{k},$
(Predictor Fitting) (4)
$\alpha^{*}=\operatorname*{arg\,min}_{\alpha\in A}f(\alpha|S^{k+1}),$
[…]
where $\text{Top}_{N}(\tilde{P}^{k})$ represents the set of the top $N$ configurations, ranked according to the task performance $\tilde{P}^{k}$. In contrast, existing methods [36] can be seen as running the above coordinate descent process for only a single iteration, which is a special case of our more general framework.
@@
By alternating between these two procedures for a predefined number of iterations, one progressively refines the predictors and guides the sampling process toward subspaces with higher-quality mixture weights, thereby increasing the average quality of the searched mixture weights. At the same time, the promising samples in $S^{k+1}$ improve the prediction accuracy of the updated predictor $f_{\theta}^{k+1}$ for high-performing configurations, allowing for more accurate assessment of the sampled configurations’ quality. Finally, one selects the best configuration predicted by the final predictor as the final data mixture weight. For implementation, the predictor can be any regression model, such as linear regression, ridge regression, decision tree regression, or a multilayer perceptron. In our experiments, we use LightGBM [37], which predicts the target value by learning an ensemble of decision trees. More implementation details could be found in Section 4.1.
@@
4 Experimental Settings
Data. For training, we use Nemotron-CC [8] and smollm-corpus [9] as the source dataset. CLIMB-clustering yields 21 super-clusters containing 800B tokens. For evaluation, we test on reasoning benchmarks: PIQA [38], ARC_C, ARC_E [39], HellaSwag [40], WinoGrande [41], and SIQA [42]. We optimize using PIQA, ARC_E, and HellaSwag validation data, then evaluate on test sets. LM-Evaluation harness [43] is used, with all datasets in a 0-shot setting except MMLU (5-shot) [44, 45].
Model. We first perform phase-1 pre-training to establish a solid foundation. Three Transformer decoder-only models (62M, 350M, 1B) are trained with next-token prediction on 10T tokens, similar to [46] (12T tokens). We use the warmup-stable-decay (WSD) learning rate schedule [47], allowing resumption in the stable stage and focusing on data mixing research in the decay stage. For proxy models, we use 62M and 350M for efficiency. For target models, we evaluate all three sizes to assess the approach across scales. For the rest of paper we use the 350M-proxy, ablations with 62M are in the Appendix A.6. Once the optimal data mixture is found, we train the target model on 40B tokens using this mixture and compare performance. Unless stated otherwise, all reported results come from this 40B continuous pre-training.
Baselines. We compare our method with (1) Random selection, and state-of-the-art data mixing methods, including (2) DoReMi [16], and (3) RegMix [36]. The details about these baselines are in Appendix A.1.1.
4.1 Implementation Details
Text embedding. We use textstella_en_400M_v5  [^2] https://huggingface.co/NovaSearch/stella_en_400M_v5, as it efficiently encodes large-scale text with excellent performance.
Embedding clustering. We adopt the classic K-means clustering algorithm from the FAISS library [48, 49], setting the initial number of clusters $K_{\text{init}}$ to 1000.
Cluster merging. We train several fasttext models [50] to evaluate the data quality across four important dimensions - overall quality, educational value, informational value, and advertisement score (1-5) - by annotating 1 million texts with Nemotron-340B [51] with a carefully designed prompt template (see Appendix A.8). Then we perform cluster-level pruning based on the fasttext scores, applying a relatively loose threshold of 3.0, which results in 240 (i.e., the value of $K_{\text{pruned}}$) clusters. Finally, we group the clusters according to a Euclidean distance threshold of 1.5, resulting in 16 clusters.
For predictor training, we use a LightGBM [37] regression model, which fits mix-performance pairs well with limited data [36]. To prevent overfitting, we set L1 and L2 regularization, early stopping, a maximum depth of four, and require at least five samples per leaf. The ablation study of the above design choices is in Section 6. Additionally, we employed a separate validation set and an early stopping mechanism, halting training after 20 rounds of no improvement.
Table 2: Comparison with state-of-the-art language models on general reasoning benchmarks. CLIMB is continuously trained on 400B tokens with the optimal data mixture. The best results are highlighted in bold.

### v2

# Nemotron-CLIMB: CLustering-based Iterative Data Mixture Bootstrapping for Language Model Pre-training

@@ Abstract
Pre-training datasets are typically collected from web content and lack inherent domain divisions. For instance, widely used datasets like Common Crawl do not include explicit domain labels, while manually curating labeled datasets such as The Pile is labor-intensive. Consequently, identifying an optimal pre-training data mixture remains a challenging problem, despite its significant benefits for pre-training performance. To address these challenges, we propose CLustering-based Iterative Data Mixture Bootstrapping (Nemotron-CLIMB), an automated framework that discovers, evaluates, and refines data mixtures in a pre-training setting. Specifically, Nemotron-CLIMB embeds and clusters large-scale datasets in a semantic space and then iteratively searches for optimal mixtures using a smaller proxy model and a predictor. This strategy enables effective domain adaptation without relying solely on curated data. When continuously trained on 400B tokens with this mixture, our 1B model exceeds the state-of-the-art Llama-3.2-1B by 2.0%. Moreover, we observe that optimizing for a specific domain (e.g., Social Sciences) yields a 5% improvement over random sampling. Finally, we introduce Nemotron-ClimbLab, a filtered 1.2T-token corpus with 20 clusters for research, and Nemotron-ClimbMix, a 400B-token compact dataset designed for efficient pre-training that delivers superior performance under an equal token budget. We analyze the final data mixture, elucidating the characteristics of an optimal data mixture. Our data is available here.
@@ 1 Introduction
Figure 1: Pre-training a 1B model on Nemotron-ClimbMix shows better scaling than training on other datasets. We measure the average performance on 12 downstream benchmarks.
Pre-training datasets for large language models (LLMs) have scaled to trillions of tokens, typically combining large-scale web crawls with smaller, high-quality domain-specific datasets. These corpora enable the development of generalist models capable of addressing diverse tasks. However, their vast scale and heterogeneity pose challenges in balancing general knowledge with domain expertise, often leading to inefficient utilization of high-value data for specialized capabilities. Recent studies emphasize the importance of the final stage of pre-training, commonly referred to as mid-training, where models are refined on targeted, high-quality data to enhance specific capabilities. For example, blakeney2024does demonstrated that emphasizing domain-specific datasets during the final pre-training phase significantly improves performance on benchmarks such as GSM8K (cobbe2021training, ) (math), MMLU (hendryckstest2021, ) (reasoning), and HumanEval (chen2021evaluating, ) (coding). Similarly, OLMo 2 (olmo20252olmo2furious, ) mixes high-quality web data with curated STEM references, synthetic math datasets, and encyclopedic content for mid-training, achieving notable gains in math reasoning tasks. These findings highlight the potential of carefully curated data mixtures in mid-training for improving domain performance.
@@
Despite the success of pre-training, optimizing data mixtures for both general and domain-specific tasks remains a challenge: (1) Large-scale datasets such as Common Crawl  [^1] https://commoncrawl.org/ offer unmatched diversity and scale but lack explicit domain labels, making it difficult to extract domain-relevant content. Filtering data often relies on general-purpose heuristics like perplexity or educational value (gunasekar2023textbooks, ), which do not necessarily capture the most informative or high-quality content for specific domains. (2) Even with curated datasets like The Pile (gao2020pile, ) with domain annotations, selecting an optimal data mixture is non-trivial due to the complex, nonlinear relationship between dataset composition and model performance. For instance, optimizing a model for coding tasks requires not just programming-related content but also complementary knowledge from mathematics, reasoning, and security.
To address these challenges, we propose CLustering-based Iterative Data Mixture Bootstrapping (Nemotron-CLIMB; CLIMB for short)—a novel framework for automating the search for optimal pre-training data mixtures. CLIMB consists of three key steps: (1) embedding and clustering large-scale datasets, (2) constructing mixture-performance pairs by sampling and pruning data mixtures and training proxy models, and (3) fitting a predictor. By treating the data mixture as input features and performance metrics as target labels, we train a regression model as a predictor. This approach enables efficient, iterative refinement of data mixtures without relying on predefined domain labels.
@@
We demonstrate the effectiveness of CLIMB by searching the optimal data mixture in general reasoning tasks first and then extending it to specific domains (e.g., STEM, social sciences, and humanities). Using the optimal data mixture discovered by CLIMB, we train 350M and 1B models on 40B tokens, both of which surpass the previously best data mixing (Doremi and RegMix) methods by a large margin. Furthermore, when trained on a larger number of tokens (400B) with this mixture, our 1B model exceeds the state-of-the-art Llama-3.2-1B by 2.0%. We observe that optimizing for a specific domain (e.g., Social Sciences) yields a 5% improvement over random sampling. Finally, based on the insights obtained from our explorations, we further apply CLIMB to two existing datasets, Nemotron-CC (su2024nemotron, ) and smollm-corpus (benallal2024smollmcorpus, ), and produce a new dataset with superior performance.
@@
    Automated Data Mixture Optimization. We propose an embedding-driven data mixing approach to automatically identify, group, mix high-quality clusters, enabling domain-specific training while removing the reliance on manually predefined domain-specific data.
@@
2 CLIMB: CLustering-based Iterative Data Mixture Bootstrapping
Our work focuses on curating training data from a massive data source in an automated fashion, specifically tailored to improve the desired tasks or domains. To ensure that the filtered dataset remains relevant to the target domain while maintaining general language modeling and reasoning capabilities, our framework simplifies the data curation process through a fully autonomous iterative bootstrapping approach, eliminating the need for manual curation and reducing labor costs. As illustrated in Figure 4, we first cluster documents from the data source in an embedding space to differentiate data across domains. Next, we iteratively optimize the mixture weights using a bootstrapping process to progressively enhance the dataset’s domain relevance. Further details of the two phases are provided in Section 2.1 and 2.2, respectively.
@@ 2.1 Data Preprocessing
Embedding clustering. We then cluster the embeddings using a suitable clustering algorithm. For instance, k-means hartigan1979algorithm can be used to group them into $K_{\text{init}}$ clusters. To ensure the clusters are as fine-grained as possible for subsequent processing, we prefer to set $K_{\text{init}}$ to a relatively large value at this stage, such as $1000$. The specific settings are detailed in Section 3.1.
@@ 2.2 Iterative Bootstrapping: Mixture Weight Search
$\min_{\alpha\in A}\,\,\ell_{val}(\alpha,\omega^{*}(\alpha))\text{s.t.}\,\,\omega^{*}(\alpha)=\operatorname*{arg\,min}_{\omega}\ell_{train}(\alpha,\omega)\text{s.t.}\sum_{i=1}^{k}\alpha_{i}=1,\quad\alpha_{i}\geq 0$ (1)
@@
(Sampling) $\tilde{P}^{k}=\{f_{k}(s)|s\in A\setminus S^{k}\},$ $S_{M}\subset\text{Top}_{N}(\tilde{P}^{k}),~S^{k+1}=S_{M}\cup S^{k},$ (3)
(Predictor Fitting) $\alpha^{*}=\operatorname*{arg\,min}_{\alpha\in A}f(\alpha|S^{k+1}),$ $\text{s.t.}\,\,f_{k+1}=\operatorname*{arg\,min}_{f_{k}\in\tilde{\mathcal{F}}}\sum_{s\in S^{k+1}}\mathcal{L}(f(s),\ell(s,\omega^{*}))$ (4)
where $\text{Top}_{N}(\tilde{P}^{k})$ represents the set of the top $N$ configurations, ranked according to the task performance $\tilde{P}^{k}$. In contrast, existing methods (liu2024regmix, ) can be seen as running the above coordinate descent process for only a single iteration, which is a special case of our more general framework.
@@
By alternating between these two procedures for a predefined number of iterations, one progressively refines the predictors and guides the sampling process toward subspaces with higher-quality mixture weights, thereby increasing the average quality of the searched mixture weights. At the same time, the promising samples in $S^{k+1}$ improve the prediction accuracy of the updated predictor $f_{\theta}^{k+1}$ for high-performing configurations, allowing for more accurate assessment of the sampled configurations’ quality. Finally, one selects the best configuration predicted by the final predictor as the final data mixture weight. For implementation, the predictor can be any regression model, such as linear regression, ridge regression, decision tree regression, or a multilayer perceptron. In our experiments, we use LightGBM (ke2017lightgbm, ), which predicts the target value by learning an ensemble of decision trees. More implementation details could be found in Section 3.1.
3 Experimental Settings
Data. For training, we use Nemotron-CC (su2024nemotron, ) and smollm-corpus (benallal2024smollmcorpus, ) as the source dataset. CLIMB-clustering yields 21 super-clusters containing 800B tokens. For evaluation, we test on reasoning benchmarks: PIQA (bisk2020piqa, ), ARC_C, ARC_E (allenai:arc, ), HellaSwag (zellers2019hellaswag, ), WinoGrande (sakaguchi2021winogrande, ), and SIQA (sap2019socialiqa, ). We optimize using PIQA, ARC_E, and HellaSwag validation data, then evaluate on test sets. LM-Evaluation harness (eval-harness, ) is used, with all datasets in a 0-shot setting except MMLU (5-shot) (allal2024SmolLM, ; smollmevalsetup, ).
Model. We first perform phase-1 pre-training to establish a solid foundation. Three Transformer decoder-only models (62M, 350M, 1B) are trained with next-token prediction on 10T tokens (a combination of DCLM (li2024datacomp, ) and TxT360 (txt360data2024, )), similar to (yang2024qwen2technicalreport, ) (12T tokens). We use the warmup-stable-decay (WSD) learning rate schedule (hu2024minicpm, ), allowing resumption in the stable stage and focusing on data mixing research in the decay stage. For proxy models, we use 62M and 350M for efficiency. For target models, we evaluate all three sizes to assess the approach across scales. For the rest of paper we use the 350M-proxy, ablations with 62M are in the Appendix D.9. Once the optimal data mixture is found, we train the target model on 40B tokens using this mixture and compare performance. Unless stated otherwise, all reported results come from this 40B continuous pre-training.
Baselines. We compare our method with (1) Random selection, and state-of-the-art data mixing methods, including (2) DoReMi (xie2024doremi, ), and (3) RegMix (liu2024regmix, ). The details about these baselines are in Appendix C.1.
3.1 Implementation Details
Text embedding. We use stella_en_400M_v5  [^2] https://huggingface.co/NovaSearch/stella_en_400M_v5, as it efficiently encodes large-scale text with excellent performance.
Embedding clustering. We adopt the classic K-means clustering algorithm from the FAISS library (johnson2019billion, ; douze2024faiss, ), setting the initial number of clusters $K_{\text{init}}$ to 1000.
Cluster merging. We train several fasttext models (joulin2016fasttext, ) to evaluate the data quality across four important dimensions - overall quality, educational value, informational value, and advertisement score (1-5) - by annotating 1 million texts with Nemotron-340B (adler2024nemotron, ) with a carefully designed prompt template (see Appendix D.11). Then we perform cluster-level pruning based on the fasttext scores, applying a relatively loose threshold of 3.0, which results in 240 (i.e., the value of $K_{\text{pruned}}$) clusters. Finally, we group the clusters according to a Euclidean distance threshold of 1.5.
For predictor training, we use a LightGBM (ke2017lightgbm, ) regression model, which fits mixture-performance pairs well with limited data (liu2024regmix, ). To prevent overfitting, we set L1 and L2 regularization, early stopping, a maximum depth of four, and require at least five samples per leaf. The ablation study of the above design choices is in Section 5. Additionally, we employed a separate validation set and an early stopping mechanism, halting training after 20 rounds of no improvement.
@@ 3.1 Implementation Details
4 Experimental Results
4.1 Comparison with Data Mixture Baselines
Table 2: Comparison with state-of-the-art language models on general reasoning benchmarks. CLIMB is continuously trained on 400B tokens with the optimal data mixture. Best results in bold.
@@ 4.1 Comparison with Data Mixture Baselines
4.2 Comparison with SOTA LMs
Using the optimal data mixture identified by our method, we further investigate the effect of scaling up. Specifically, we used the same data mixture to train on 400B tokens and then compared the resulting model against state-of-the-art baselines. As shown in Table 2, CLIMB achieves the best performance among all sub-500M and sub-1.2B models. For example, when comparing models of similar scales (around 1B parameters), CLIMB consistently outperforms other baselines — including Llama-3.2 and AMD-OLMo — across the majority of the general reasoning benchmarks. In particular, it achieves the highest overall average score, surpassing the next-best model (i.e., Llama-3.2) by a noticeable margin (2.0%). Moreover, we introduced additional benchmarks (e.g., mmlu, gpqa, obqa, boolq, and race), and our model is consistently better than baseline models, which demonstrates that our method exhibits excellent generalization performance.
5 Analysis